This is an OCR A Gateway spec A level Biology cheat sheet for Chapter 20 module 6. Specification reference: 6.1.2
|
Genotype
|
Genetics makeup of an organisms - the alleles present.
|
|
Phenotype
|
The expression of genes (also affected by the environment).
|
|
Homozygous
|
A pair of homologous chromosomes carrying same allele for a gene (e.g. AA, aa).)
|
|
Heterozygous
|
A pair of homologous chromosomes carrying different alleles for a gene (e.g. Aa).
|
|
Recessive allele
|
Allele only expressed in the absence of dominant alleles
|
|
Dominant allele
|
Allele always expressed in phenotype.
|
|
Codominant
|
Multiple alleles are equally dominant and expressed in the phenotype.
|
|
Sex-linkage
|
A gene with a locus in the X chromosome.
|
|
Autosomal linkage
|
Genes in the same chromosome (not sex chromosome). An autosomal chromosome is any chromosome other than sex chromosomes.
|
|
Epistasis
|
When one gene modifies the expression of a different gene.
|
|
Monohybrid
|
Genetics inheritance cross of trait determined by one gene.
|
|
Dihybrid
|
Genetics inheritance cross of trait determined by two gene.
|
|
Gene pool
|
All alleles of all genes within population.
|
|
Allele frequency
|
Proportion of an allele in a gene pool.
|
|
Type of inheritance
|
Coding
|
|
Monohybrid
|
Single letter capital / lower case e.g. A, a
|
|
Codominance
|
Geneallele e.g. CWCR, IAIB
|
|
Multiple alleles for one gene (more than two)
|
Geneallele e.g. IAIO
|
|
Sex linkage
|
Chromosomeallele e.g. XRXr, XrY
|
|
Autosomal linkage and epistasis
|
Single letters e.g. Aa Bb
|
|
Worked example: Cystic fibrosis is caused by a recessive allele, what is the probability that two carrier parents will have a child with cystic fibrosis?
|
||
|
Parent genotypes ("both carriers"): Ff x Ff
|
||
|
Each gamete will carry either the F or f allele and can fuse with gametes containing either F or f alleles too because parents are heterozygous.
|
||
|
F
|
f
|
|
|
F
|
FF (no CF)
|
Ff (no CF)
|
|
f
|
Ff (no CF)
|
ff (CF)
|
|
Therefore probability of child with cystic fibrosis = 1/4 = 25%.
|
||
|
Worked example: A flower can be three colours: white, red or pink. The alleles for white and red are codominant to form pink flowers. If two pink flowers reproduce, what is the probability of forming a red offspring flower?
|
||
|
Pink flowers must have the alleles both for red and white as they are codominant to produce pink. They are therefore heterozygous.
|
||
|
Parent genotype: CRCW x CRCW
|
||
|
CR
|
CW
|
|
|
CR
|
CRCR (red)
|
CRCW (pink)
|
|
CW
|
CRCW (pink)
|
CWCW (white)
|
|
Therefore probability of forming a red offspring is 1/4 = 25%.
|
||
|
Key example: Blood groups
|
||
|
There are three alleles for blood group: A, B and O, represented as IA, IB and IO.
|
||
|
Both IA and IB are codominant to form phenotype group AB, and IO is recessive.
|
||
|
Worked example: parents with blood group AB and O reproduce. WHat so the probability that they will produce an offspring with blood group A?
|
||
|
Parent genotype: IAIB x IOIO
|
||
|
IA
|
IB
|
|
|
IO
|
IAIO (group A)
|
IBIO (group B)
|
|
IO
|
IAIO (group A)
|
IBIO (group B)
|
|
Therfore probability is 2/4 = 50%.
|
||
|
Sex linked alleles only occur in X chromosomes as Y chromosomes contain less genetic information (too small to carry more). This makes males more likely to carry recessive sex-linked disorders as their homologous Y chromosome cannot contain the dominant allele.
|
||
|
Worked example: Colour blindness is caused by a recessive sex-linked allele only found in the X chromosome. If a colour blind male reproduces with a heterozygous female, what is the probability that they would have a colour blind child?
|
||
|
Parent genotype: XBXb x XbY
|
||
|
XB
|
Xb
|
|
|
Xb
|
XBXb
|
XbXb
|
|
Y
|
XBY
|
XbY
|
|
The probability of a colour blind child is therefore 2/4 = 50%.
|
||
|
One gene affects the expression of another, therefore multiple genes will be at play here.
|
||||
|
(Common) worked example: Labradors have two genes which will affect their fur colour. The first gene controls which colour is expressed. If allele B is expressed, the dog will be black. If allele b is expressed, the dog will be brown. The second gene codes for pigment production. If allele E is expressed, pigment will be produced. If allele e is expressed, no pigment will be produced and the dog will be yellow. Parents heterozygous for both genes reproduce.
|
||||
|
Parent genotype: Bb Ee x Bb Ee
|
||||
|
Parent possible gametes: BE, Be, bE, be x BE, Be, bE, be
|
||||
|
BE
|
Be
|
bE
|
be
|
|
|
BE
|
BBEE
|
BBEe
|
BbEE
|
BbEe
|
|
Be
|
BBEe
|
BBee
|
BbEe
|
Bbee
|
|
bE
|
BbEe
|
BbEe
|
bbEE
|
bbEe
|
|
be
|
BbEe
|
Bbee
|
bbEe
|
bbee
|
|
Offspring phenotypes:
|
||||
|
Black : Brown : Yellow
|
||||
|
9:3:4
|
||||
|
Two genes considered at the same time.
|
||||
|
Key example: Mendel's peas
|
||||
|
Peas can be two different colours, yellow or green. They can also be either round or wrinkled. Roundness is a dominant allele (R) and yellow colour is also dominant (Y). Show the proportions of phenotypes for the offsprings of two heterozygous peas.
|
||||
|
Parent genotype: RrYy x RrYy
|
||||
|
Gametes: RY, Ry, rY, ry x RY, Ry, rY, ry
|
||||
|
RY
|
Ry
|
rY
|
ry
|
|
|
RY
|
RRYY
|
RRYy
|
RrYY
|
RrYy
|
|
Ry
|
RRYy
|
RRyy
|
RrYy
|
Rryy
|
|
rY
|
RrYy
|
RrYy
|
rrYY
|
rrYy
|
|
ry
|
RrYy
|
Rryy
|
rrYy
|
rryy
|
|
Offspring phenotypes:
|
||||
|
Round and yellow : Wrinkled and yellow : Round and green : Wrinkled and green
|
||||
|
9:3:3:1
|
||||
|
This ratio will always be present in dihybrid heterozygous crosses.
|
||||
|
When genes are linked, it means they occur on the same chromosome.
For example, R and Y are on one chromosome and r and you on the other. |
||
|
This means it is not possible to get all the gametes predicted in the previous box.
R and y cannot form a gamete because R and Y are always inherited together. |
||
|
In the example we used, the only possible gametes for heterozygous parents if the genes are linked are:
RY, ry x RY, ry |
||
|
RY
|
ry
|
|
|
RY
|
RRYY
|
RrYy
|
|
ry
|
RrYy
|
rryy
|
|
The ratio therefore changes from 9:3:3:1 to 3:1 (in this example.)
|
||
|
Crossing over (meiosis)
|
||
|
It is, however, possible for observations to show more than just two phenotypes, or unexpected proportions of these phenotypes.
|
||
|
This is because of crossing over in meiosis. Homologous chromosomes can overlap and swap parts of non-siter chromatids. This can form new gametes.
In our example, part of the chromosome containing the Y and another chromatid containing the y allele could swap, creating the gametes Ry and rY. These are called recombinant genes. |
||
|
Genes are less likely to split if they are close together, because there is less space for a chiasmata to form between them so they are less likely to be separated and will be inherited together (linked).
|
||
|
Statistical test used to calculate whether what we expect is different from what we actually observe.
|
||
|
(O-E)2/E
|
Degrees of freedom = number of categories - 1
|
|
|
Worked example: Corn can be yellow (Y) or purple (y). Heterozygous cross - Mendelian genetics expected:
|
||
|
Y
|
y
|
|
|
Y
|
YY
|
Yy
|
|
y
|
Yy
|
yy
|
|
yellow : purple = 3:1
|
||
|
A student observed 21 yellow and 13 purple kernels. Is this significantly different from expectation?
|
||
|
Null hypothesis: no significant difference between expected and observed colour of corn.
|
||
|
1. Make expected a proportion
|
21+13 = 34
3+1 = 4 34/4 = 8.5 3x8.5 = 25.5 1x8.5 = 8.5 |
|
|
2. Observed - Expected
|
21-25.5 = -4.5
13-8.5 = 4.5 |
|
|
3. Square answers
|
-4.52 = 20.25
4.52 = 20.25 |
|
|
4. Divide by expected value
|
20.25/25.5 = 0.794
20.25/8.5 = 2.382 |
|
|
5. Add up values
|
Chi squared = 3.176
|
|
|
6. Calculate degrees of freedom and compare calculated value to critical value table
|
2-1 = 1
critical value = 3.841 |
|
|
Because our calculated value is lower than the critical value, there is more than 5% probability that any differences is due to chance. This means results are not significantly different and we accept the null hypothesis.
|
||
|
So the colour observed matched the Mendelian genetics we originally calculated.
|
||
|
Mathematical model used to determine the frequency of alleles.
|
|
|
p2 + 2pq + q2 = 1
|
|
|
p2 = frequency of homozygous dominant genotype (AA)
2pq = frequency of heterozygous genotype (Aa) q2 = frequency of homozygous recessive genotype (aa) |
|
|
p + q = 1
|
|
|
p = frequency of dominant allele
q = frequency of |
|
|
Worked example: Haemochromatosis is caused by a recessive allele. In one country, every 1 in 400 people have haemochromatosis. What percentage of the population is a carrier for the haemochromatosis recessive gene?
|
|
|
The data given to us is the proportion of homozygous recessive individuals, so q2.
Therefore, q2 = 1/400 = 0.0025. |
|
|
1. Use the information given to work out other values.
If we know q2, then we can find q then p. |
√q2 = q
√0.0025 = q = 0.05 p+q = 1, therefore p + 0.05 = 1 1-0.05 = p = 0.95 |
|
2. Substitute for answer
|
Carriers = heterozygous, so 2pq
2x0.95x0.05 = 0.095 0.095x100 = 9.5% |
|
Creation of new species. A group gets reproductively isolated from the population (cannot breed together) and develop differences in gene pools.
|
|
|
Allopatric speciation
|
Become geographically isolated (e.g. new mountain range). Evolve separately through natural selection, accumulate mutations.
|
|
Sympatric speciation
|
Occupy same area but behaviours change so don't breed together.
|
|
Change in allele frequency in population. Impacts are greater in smaller populations.
|
|
|
Factors that can lead to genetic drift are:
|
|
|
Genetic bottlenecks
|
Event (e.g. natural disaster, overfishing...) kills off most of a population, leaving a few survivors behind (small gene pool).
|
|
Founder effect
|
A few individuals (small gene pool) first colonise an area, isolated from original population. Can even make rare homozygotic recessive phenotypes more frequent.
|
|
Natural selection
|
Evolution. Variety in phenotypes due to genetic bad environmental factors.
|
|
Artificial selection
|
Humans select desirable features and breed those individuals together.
|
|
Ethical concern: due to selected features, some animals (e.g. pugs) will have medical issues.
|
|
Gene banks store DNA from plants or animals.
|
|
Selective breeding often involves inbreeding, so gene banks can be used to reduce this and increase genetic diversity of a species.
|
No comments yet.
Cheatography n. The art of mapping complex subjects into clear, instantly navigable reference guides.